# Scaling Up Active Testing to Large Language Models

### v1

@@ Abstract
Active testing enables label-efficient evaluation of models through careful data acquisition. However, its significant computational costs have previously undermined its use for large models. We show how it can be successfully scaled up to the evaluation of large language models (LLMs). In particular we show that the surrogate model used to guide data acquisition can be constructed cheaply using in-context learning, does not require updating within an active-testing loop, and can be smaller than the target model. We even find we can make good data-acquisition decisions without computing predictions with the target model and further introduce a single-run error estimator to asses how well active testing is working on the fly. We find that our approach is able to more effectively evaluate LLM performance with less data than current standard practices.
@@ 1 Introduction
Evaluating frontier models is becoming more expensive as they become more sophisticated [5, 25] and currently represents a critical challenge in the field [27, 36, 39]. Moreover, with new models arriving in quick succession, and ever-present scope for data to leak from evaluations to training [9], the evaluation problem is a dynamic one: it requires ongoing, adaptive gathering of new evaluation data.
@@
Active testing [16, 17] is an attractive solution to this problem. Motivated by the observation that some labels are more informative than others about the behaviour of a target model, it involves carefully, and typically adaptively, deciding which test inputs to acquire labels for. The foundation for this is a surrogate model of the test-time label distribution, which is typically updated on the labels acquired during testing. Making a data-acquisition decision requires computing predictions with the surrogate model and often also the target model. The computational cost of this process is often justifiable in light of high labelling costs—but it limits the scope of active testing’s applicability.
@@
In this work, we show how active testing can be scaled up to LLMs, thereby allowing them to be more efficiently evaluated with less labelled test data. To this end, we first highlight three potential computational bottlenecks in the active testing pipeline: updating the surrogate model as labels are acquired, computing predictions with the surrogate model, and computing predictions with the target model. The first of these is particularly prominent: it has historically involved repeatedly running gradient-based training on the acquired test data within an active-testing loop.
@@
Figure 1: Our proposed active-testing approach enables low-error estimates of the risk (expected predictive test loss) of a large language model (here the 70B version of Llama 2) on four text-classification problems (SST-2, FPB, HS and Subj) while only using a small label budget. We compare uniform-random testing against active testing (LURE), using either a 7B or 70B surrogate model with in-context learning to guide data acquisition.
@@
We then identify simple and surprisingly effective ways to overcome each of these key bottlenecks. First, we strip back the iterative training of the underlying surrogate-model to a single step of in-context learning [4] on a small amount of initial test data, removing the need for repeated gradient-based training. Second, we show it is possible to use a surrogate model that is smaller than the target model, reducing the cost of surrogate-model predictions. Third, we show that one can even forgo computing predictions with the target model and rely solely on the surrogate model for data acquisition. Together these changes allow us to scale up active testing to LLMs.
Empirically, our approach substantially improves over the standard practice of acquiring data at random (Figure˜1), typically reducing estimation error by 25% to 50%, and sometimes up to 80%.
@@
To further improve the real-world applicability of active testing, we also address the challenge of judging how well it is working on a given problem. Practical deployment typically involves running active testing only once, without knowledge of the true risk or the ability to rerun the approach multiple times. To help gauge performance in this setting, we derive a bootstrap estimator [7] of the risk-estimation error. Experimentally our approximate 95% confidence intervals contained the true risk-estimation error 88% of the time, demonstrating their efficacy as a diagnostic tool.
In summary, our work allows for LLMs to be evaluated more accurately and with less labelled test data than current best practices by actively selecting data to label at test-time.
@@ 2 Background
Our aim is to evaluate a fixed target model, $f$, mapping inputs $x\in\mathcal{X}$ to outputs $y\in\mathcal{Y}$. We formalise evaluation as estimating the frequentist risk of $f$ [3], i.e. the expected loss $R=\mathbb{E}_{p_{\mathrm{eval}}(x,y)}\!\left[\ell(f(x),y)\right]$, where $p_{\mathrm{eval}}$ represents a reference system used as a source of ground truth, and $\ell$ denotes a loss function. As in past work on active testing [16, 17], we consider a pool-based setting [20] where we have access to a pool of $N$ unlabelled test inputs, $\mathcal{D}_{\mathrm{pool}}=\{x_{i}\}_{i=1}^{N}$, but acquiring a label, $y\sim p_{\mathrm{eval}}(y|x)$, for any given $x$ is costly, so we can only afford to acquire $M<N$ labels.
@@
Uniform-random sampling A straightforward way to estimate $R$ is to sample $M$ inputs uniformly at random from the pool and compute $\hat{R}_{\mathrm{unif}}=\frac{1}{M}\sum_{m=1}^{M}\ell(f(x_{i_{m}}),y_{i_{m}})$, where $i_{1:M}\sim\mathrm{Uniform}(\{1,2,\ldots,N\},M)$ are samples drawn without replacement. If the pool was constructed by sampling $x_{i}\sim p_{\mathrm{eval}}(x)$, then $\hat{R}_{\mathrm{unif}}$ is known as the subsample empirical risk and is an unbiased estimator of $R$. However, it will typically have high variance for small values of $M$.
@@
Sampling-based active testing Active testing deals with how to more carefully select the $M$ inputs for labelling to produce a more accurate estimate of $R$. In general this can be done in either a sampling-based or interpolation-based manner. The sampling-based approach uses Monte Carlo estimators of $R$, similar to $R_{\mathrm{unif}}$ but with acquisition indices sampled from a non-uniform distribution over the pool indices. Namely, at acquisition step $m$ we sample pool input $x_{i}$ with probability proportional to $a_{m}(x_{i})\in\mathbb{R}^{+}$, $q_{m}(i)=\frac{a_{m}(x_{i})}{\sum_{x_{j}\in\mathcal{D}_{\mathrm{pool}}}a_{m}(x_{j})}$, where $a_{m}$ is an acquisition function that measures a notion of the expected value of acquiring a label for a given input (the $m$ subscript denotes that it can depend on all the data available at step $m$, including the target model’s training data).
@@
A naive estimator of the risk is $\hat{R}_{\mathrm{naive}}=\frac{1}{M}\sum_{m=1}^{M}\ell(f(x_{i_{m}}),y_{i_{m}})$, where $i_{m}\sim q_{m}(i)$. Farquhar et al, [8] showed that $\hat{R}_{\mathrm{naive}}$ is biased and introduced the levelled unbiased risk estimator (LURE):
$\hat{R}_{\mathrm{LURE}}=\frac{1}{M}\sum_{m=1}^{M}v_{m}\ell(f(x_{i_{m}}),y_{i_{m}}),\quad v_{m}=1+\frac{N-M}{N-m}\left(\frac{1}{(N-m+1)q_{m}(i_{m})}-1\right),$ (1)
where $i_{m}\sim q_{m}(i)$. Given both $\hat{R}_{\mathrm{unif}}$ and $\hat{R}_{\mathrm{LURE}}$ are unbiased, any advantage that $\hat{R}_{\mathrm{LURE}}$ brings comes from variance reduction through a well-designed acquisition function, $a_{m}$.
@@
The optimal acquisition function is $a^{*}_{m}(x)=\mathbb{E}_{p_{\mathrm{eval}}(y|x)}\!\left[\ell(f(x),y)\right]$, i.e. the expected loss under $p_{\mathrm{eval}}(y|x)$. Since this is unknown, Kossen et al, [16] proposed approximating $p_{\mathrm{eval}}(y|x)$ with a surrogate model, $\pi_{m}(y|x)$, which is typically trained on the acquired test labels along with the target model’s training data. This then allows us to define a practical acquisition function:
@@
0: Target model, $f$; loss function, $\ell$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; budget, $M$
@@
8:  Compute $\ell(f(x_{i_{m}}),y_{i_{m}})$ and $v_{m}$
9:  Set $\mathcal{D}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup\{(x_{i_{m}},y_{i_{m}})\}$
10:end for
11: Compute $\hat{R}_{\mathrm{LURE}}$
11: Risk estimate, $\hat{R}_{\mathrm{LURE}}$
@@
0: Target model, $f$; loss function, $\ell$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; budget, $M$; acquisition function, $a$
@@
2: Set $\mathcal{D}_{\mathrm{test}}=\emptyset$.
@@
5:  Compute $a(x_{j},\pi_{m})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
6:  Select $i_{m}=\arg\max_{j}a(x_{j},\pi_{m})$
@@
Interpolation-based active testing Kossen et al, [17] introduced an alternative approach to active testing that, like the sampling-based approach, uses a surrogate model to guide data acquisition, but also uses it to estimate the risk. Their active surrogate estimator (ASE) is
@@
Here the risk is not directly estimated using labels acquired from $y\sim p_{\mathrm{eval}}(y|x)$ but instead using labels simulated from the surrogate model. The goal of data acquisition is then to improve the surrogate model in a way that leads to a more accurate estimate of the risk. Kossen et al, [17] approached this using an acquisition function called the expected weighted disagreement (XWED).
@@ 3 Scaling Up Active Testing
With an understanding of standard approaches, we can now address barriers to scaling up active testing to LLMs. We focus on three key computational bottlenecks: training the surrogate model, computing predictions with the surrogate model, and computing predictions with the target model. Our aim is to reduce the cost of these steps while maintaining the efficacy of active testing.
@@ 3.1 Surrogate-model training
To avoid this expense, we suggest a stripped-back approach: construct the surrogate model using in-context learning [4] on a small amount of randomly acquired data, and fix the surrogate model thereafter. This reduces the cost of surrogate-model training to essentially the minimum possible (other than no training at all). This design decision has important implications for the relative merits of sampling-based ($\hat{R}_{\mathrm{LURE}}$) and interpolation-based ($\hat{R}_{\mathrm{ASE}}$) active testing. We are choosing to use a relatively crude surrogate model to reduce computational cost. This affects $\hat{R}_{\mathrm{LURE}}$ solely through the data-acquisition step (and related corrective weights $v_{m}$), whereas it affects $\hat{R}_{\mathrm{ASE}}$ not only through data acquisition but also much more directly through the expectation over $\pi_{m}(y|x)$ in Equation˜3. Due to this difference in sensitivity between the approaches, we choose $\hat{R}_{\mathrm{LURE}}$ as our default risk estimator for this choice of surrogate model, and demonstrate in Section˜5 that this is well-justified empirically.
@@ 3.2 Surrogate-model predictions
Another cost is that of computing predictions with the surrogate model. In sampling-based active testing these predictions arise when computing the acquisition distribution, $q_{m}(i)$, and in interpolation-based active testing they additionally arise when computing the risk estimate, $\hat{R}_{\mathrm{ASE}}$. Our choice to use a fixed surrogate model, $\pi_{0}$, automatically leads to a significant reduction in how many of these predictions are required: we only need to compute $\pi_{0}(y|x_{j})$ once for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$. On top of this we can make each individual forward pass cheaper by using a smaller surrogate model.
@@ 3.3 Target-model predictions
The third cost we aim to reduce is that of computing predictions with the target model. The standard versions of both sampling-based and interpolation-based active testing require computing $f(x_{j})$ for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$, which is especially expensive for large target models and pool sets.
@@
where $\psi$ denotes an operation that accounts for the fact that the first argument of $\ell$ might not be a probability distribution (e.g., if $\ell(\hat{y},y)=\|\hat{y}-y\|_{2}^{2}$, then we require something like $\psi[\pi_{m}(\cdot|x)]=\mathbb{E}_{\hat{y}\sim\pi_{m}(\cdot|x)}\!\left[\hat{y}\right$]). This acquisition function can be seen as an approximation to $a_{m}(x)$ in Equation˜2. If $\ell(\hat{p},y)=-\log\hat{p}(y)$ then $\hat{a}_{m}(x)=\mathrm{H}\!\left[\pi_{m}(y|x)\right]$, the predictive entropy of the surrogate model.
@@ 3.4 Active testing for dataset curation
The use of the surrogate model to approximate the target model during data acquisition also makes it possible to use active testing for dataset curation, i.e. when all the inputs in the pool set are already labelled and the goal is to reduce the number of target-model predictions required for evaluation [23, 28, 29, 35]. This can be viewed as a variation on standard active testing where here we want to use the surrogate model to select $M<N$ input-label pairs with which to evaluate the target model. The acquisition function in this case can depend on the labels as well as the inputs:

### v2

@@ Abstract
Active testing enables label-efficient evaluation of predictive models through careful data acquisition, but it can pose a significant computational cost. We identify cost-saving measures that enable active testing to be scaled up to large language models (LLMs). In particular we show that the surrogate model used to guide data acquisition can be constructed cheaply using in-context learning, does not require updating within an active-testing loop, and can be smaller than the target model. We even find we can make good data-acquisition decisions without making predictions with the target model. As a result we are able to achieve much more accurate evaluations of LLM performance relative to using randomly acquired data. We additionally introduce a bootstrap estimator of evaluation error, which we show to be a useful indicator of how well active testing is working within a single run.
@@ 1 Introduction
Evaluating frontier models is becoming more expensive as they become more sophisticated (burden2024evaluating; openai2023gpt4). At the same time, with new models arriving in quick succession, and ever-present scope for data to leak from evaluations to training (ganguli2023challenges), the evaluation problem is a dynamic one: it requires ongoing, adaptive gathering of new evaluation data.
@@
Active testing (kossen2021active; kossen2022active) is an attractive solution to this problem. Motivated by the observation that some labels are more informative than others about the behaviour of a target model, it involves carefully deciding which test inputs to acquire labels for. The foundation for data acquisition is a surrogate model of the test-time label distribution, which is typically updated on the labels acquired during testing. Making a data-acquisition decision requires making predictions with the surrogate model and often also the target model. The computational cost of this process is often justifiable in light of high labelling costs, but it limits the scope of active testing’s applicability.
@@
In this work we show how active testing can be scaled up to large language models (LLMs), allowing us to more effectively evaluate them with limited labelling budgets. We begin by highlighting three computational bottlenecks in the predominant approach to active testing: updating the surrogate model as labels are acquired, making predictions with the surrogate model, and making predictions with the target model. The first of these is particularly important: it has traditionally involved repeatedly running gradient-based training on the acquired test data within an active-testing loop.
@@
Figure 1: Our proposed active-testing approach enables low-error estimates of the risk (expected predictive loss) of a large language model (here the 70B version of Llama 2) on four text-classification problems (SST-2, FPB, HS and Subj) while only using a small label budget. We compare uniform-random testing against active testing (LURE), using either a 7B or 70B surrogate model with in-context learning to guide data acquisition.
@@
We then identify simple and surprisingly effective ways to address each of these key bottlenecks. First, we strip back the training of the surrogate model to a single step of in-context learning (brown2020language) on a small amount of initial test data, removing the need for repeated gradient-based training. Second, we show it is possible to use a surrogate model that is smaller than the target model, reducing the cost of surrogate-model predictions. Third, we show we can even forgo making any predictions with the target model, relying solely on the surrogate model for data acquisition. Together these changes make active testing dramatically cheaper, enabling it to be scaled up to LLMs.
Empirically we find our approach can substantially improve over the common practice of acquiring data uniformly at random (Figure 1). More specifically it can estimate the risk (expected predictive loss over test examples) of LLMs with an estimation error typically 25% to 50%—and sometimes up to 80%—lower than that obtained through uniform-random data acquisition. Our work therefore represents significant progress in accurately and dynamically evaluating LLMs, as well as in more generally increasing the scope for active testing to be used in practical settings.
@@
To further improve the real-world applicability of active testing, we address the challenge of judging how well it is working on a given problem. In research we can assess active testing by running it multiple times and comparing its risk estimates to a known true risk. But practical deployment involves running active testing only once, without knowledge of the true risk, making it hard to gauge active-testing performance. To help address this we derive a bootstrap estimator (efron1979bootstrap) of the risk-estimation error. In experiments we find our approximate 95% confidence intervals contain the true risk-estimation error 88% of the time, suggesting our estimator can be a useful diagnostic tool.
@@ 2 Background
Our aim is to evaluate a fixed target model, $f$, mapping inputs $x\in\mathcal{X}$ to outputs $y\in\mathcal{Y}$. We formalise evaluation as estimating a form of frequentist risk (berger1985statistical), namely an expected predictive loss of the form $R=\mathbb{E}_{p_{\mathrm{eval}}(x,y)}\!\left[\ell(f(x),y)\right]$, where $p_{\mathrm{eval}}$ represents a reference system used as a source of ground truth, and $\ell$ denotes a loss function that represents the consequences of predictive errors. As in past work (kossen2021active; kossen2022active), we consider a pool-based setting (lewis1994sequential) where we have access to a pool of $N$ unlabelled test inputs, $\mathcal{D}_{\mathrm{pool}}=\{x_{i}\}_{i=1}^{N}$, but acquiring a label, $y\sim p_{\mathrm{eval}}(y|x)$, for any given $x$ is costly, so we can only afford to acquire $M<N$ labels.
2.1 Uniform-random sampling
A simple estimator of $R$ uses uniform-random samples from the pool:
$\hat{R}_{\mathrm{unif}}=\frac{1}{M}\sum_{m=1}^{M}\ell(f(x_{i_{m}}),y_{i_{m}}),$
@@ 2.1 Uniform-random sampling
where $i_{1:M}\sim\mathrm{Uniform}(\{1,2,\ldots,N\},M)$ are indices sampled without replacement. If the pool was constructed by sampling $x_{i}\sim p_{\mathrm{eval}}(x)$ then $\hat{R}_{\mathrm{unif}}$ is known as the subsample empirical risk and is an unbiased estimator of $R$. However, it will typically have high variance for small values of $M$.
2.2 Sampling-based active testing
@@
Active testing deals with how to more carefully select the $M$ inputs for labelling to produce a more accurate estimate of $R$. This can be achieved through either a sampling-based approach or an interpolation-based approach. In sampling-based active testing we use Monte Carlo estimators of $R$, similar to $R_{\mathrm{unif}}$ but with pool indices sampled from a non-uniform distribution. At acquisition step $m$ we sample pool input $x_{i}$ with probability proportional to $a_{m}(x_{i})\in\mathbb{R}^{+}$ where $a_{m}$ is an acquisition function that measures a notion of how useful we expect the label for a given input to be (the $m$ subscript denotes that it can depend on all the data available at step $m$, including the target model’s training data). That is, we sample $i_{m}\sim q_{m}(i)$ where $q_{m}(i)=a_{m}(x_{i})/\sum_{x_{j}\in\mathcal{D}_{\mathrm{pool}}}a_{m}(x_{j})$.
@@
farquhar2021statistical showed that naive Monte Carlo with $i_{m}\sim q_{m}(i)$ is biased and, to address this, introduced the levelled unbiased risk estimator (LURE):
$\hat{R}_{\mathrm{LURE}}=\frac{1}{M}\sum_{m=1}^{M}v_{m}\ell(f(x_{i_{m}}),y_{i_{m}}),\quad v_{m}=1+\frac{N-M}{N-m}\left(\frac{1}{(N-m+1)q_{m}(i_{m})}-1\right).$ (1)
@@
Given both $\hat{R}_{\mathrm{unif}}$ and $\hat{R}_{\mathrm{LURE}}$ are unbiased, any advantage that $\hat{R}_{\mathrm{LURE}}$ brings comes from variance reduction through a well-designed acquisition function, $a_{m}$. The optimal acquisition function is $a^{*}_{m}(x)=\mathbb{E}_{p_{\mathrm{eval}}(y|x)}\!\left[\ell(f(x),y)\right]$, the expected loss under $p_{\mathrm{eval}}(y|x)$. Since this is unknown, kossen2021active proposed approximating $p_{\mathrm{eval}}(y|x)$ with a surrogate model, $\pi_{m}(y|x)$, which is typically trained on the acquired test labels along with the target model’s training data. This surrogate model then allows us to define a practical acquisition function:
@@
0: Target model, $f$; loss function, $\ell$; acquisition function, $a$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; label budget, $M$
@@
9:end for
10: Compute $\hat{R}_{\mathrm{LURE}}$
10: Risk estimate, $\hat{R}_{\mathrm{LURE}}$
@@
0: Target model, $f$; loss function, $\ell$; acquisition function, $a$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; label budget, $M$
@@
2: Set $\mathcal{D}_{\mathrm{test}}=\emptyset$
@@
5:  Compute $a_{m}(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
6:  Select $i_{m}=\arg\max_{j}a_{m}(x_{j})$
@@
2.3 Interpolation-based active testing
kossen2022active introduced an alternative approach to active testing that, like the sampling-based approach, uses a surrogate model to guide data acquisition, but also uses it to estimate the risk. Their active surrogate estimator (ASE) is
@@ 2.3 Interpolation-based active testing
Here the risk is not directly estimated using labels acquired from $y\sim p_{\mathrm{eval}}(y|x)$ but instead using labels simulated from the surrogate model. The goal of data acquisition is then to improve the surrogate model in a way that leads to a more accurate estimate of the risk. kossen2022active approached this using an acquisition function called the expected weighted disagreement (XWED).
@@ 3 Scaling up active testing
With an understanding of standard approaches to active testing, we can now identify and address barriers to scaling up to LLMs. We focus on three key computational bottlenecks: training the surrogate model, making predictions with the surrogate model, and making predictions with the target model. Our aim is to reduce the cost of these steps while maintaining the efficacy of active testing.
@@ 3.1 Surrogate-model training
To avoid this expense we suggest a stripped-back approach: construct the surrogate model using in-context learning (brown2020language) on a small amount of randomly acquired test data, then keep it fixed. This reduces the cost of surrogate-model training to essentially the minimum possible.
This design decision has important implications for the relative merits of sampling-based and interpolation-based active testing. We are using a relatively crude surrogate model to reduce computational cost. This affects the sampling-based estimator solely through the data-acquisition step (and related corrective weights, $v_{m}$), whereas it affects the interpolation-based estimator not only through data acquisition but also much more directly through the expectation over $\pi_{m}(y|x)$ in Equation 3. Due to this difference in sensitivity between the approaches, we default to the sampling-based estimator for this choice of surrogate model, and demonstrate in Section 5 that this is well-justified empirically.
@@ 3.2 Surrogate-model predictions
Another cost is that of making predictions with the surrogate model. In sampling-based active testing these predictions arise when computing the acquisition distribution, $q_{m}(i)$, and in interpolation-based active testing they additionally arise when computing the risk estimate.
Our choice to use a fixed surrogate model, $\pi_{0}$, automatically leads to a significant reduction in how many of these predictions are required: we only need to compute $\pi_{0}(y|x_{j})$ once for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$. On top of this we can make each individual forward pass cheaper by using a smaller surrogate model.
@@ 3.3 Target-model predictions
The third cost we aim to reduce is that of making predictions with the target model. The standard versions of both sampling-based and interpolation-based active testing require computing $f(x_{j})$ for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$, which is especially expensive for large target models and pool sets.
@@
where $\psi$ denotes an operation that accounts for the fact that the first argument of $\ell$ might not be a probability distribution (e.g., if $\ell(\hat{y},y)=\|\hat{y}-y\|_{2}^{2}$, then we require something like $\psi[\pi_{m}(\cdot|x)]=\mathbb{E}_{\hat{y}\sim\pi_{m}(\cdot|x)}\!\left[\hat{y}\right$]). This acquisition function can be seen as an approximation to $a_{m}(x)$ in Equation 2. If $\ell(\hat{p},y)=-\log\hat{p}(y)$ then $\hat{a}_{m}(x)=\mathrm{H}\!\left[\pi_{m}(y|x)\right]$, the predictive entropy of the surrogate model.
@@ 3.4 Active testing for dataset curation
The use of the surrogate model to approximate the target model during data acquisition also makes it possible to use active testing for dataset curation, where all the inputs in the pool set are already labelled and the goal is to reduce the number of target-model predictions required for evaluation (maynez2023benchmarking; polo2024tinybenchmarks; saranathan2024dele; vivek2024anchor). This can be viewed as a variation on standard active testing where here we want to use the surrogate model to select $M<N$ input-label pairs with which to evaluate the target model. The acquisition function in this case can depend on the labels as well as the inputs: